OpenStudy (anonymous):

Use an addition or subtraction formula to find the exact value of the expression of sin(-705)

OpenStudy (anonymous):

\[\sin((-705)°)\] \[\frac{ \sqrt{3} - 1 }{ 2\sqrt{2} }\] \[\approx 0.25881\]

OpenStudy (anonymous):

first off \(-705\) is co terminal with \(15\)

OpenStudy (anonymous):

as \(-705+360+360=15\)

OpenStudy (anonymous):

so you are asked to compute \(\sin(15)\) which you can use the half angle formula for, since \(15=\frac{30}{2}\)